i need the slope of the line 3x+5y=2
Answer by oberobic(2304)

3x +5y = 2
.
Use your knowledge to get the equation into slope-intercept form.
y = mx + b
m = slope
(0,b) = y-intercept
.
5y = -3x +2
y = 3/5*x + 2/5
.
m = 3/5 = slope
.
Done.